A company has cost of goods available for sale of $250,000, sales of \$287,000 , and a gross profit percentage of 30 percent. Using the gross profit method, what is the ending inventory? a. $49.100 b. $50,000 c. $113,000 d. \$163,900
Answer #1

gross profit%=gross profit/Net sales

gross profit=(287,000*30%)=$86100

gross profit=Sales-Cost of goods sold

Cost of goods sold=287,000-86100

=$200900

Cost of goods sold=cost of goods available for sale-ending inventory
ending inventory=250,000-200900

=$49100.